OpenStudy (anonymous):

There are three cabinets, A, B, and C,each of which has 2 drawers. Each drawer has 1 coin: A has 2 gold coins, B has 2 silver coins and C has 1 silver and one gold coin. A cabinet is chosen at random, one drawer is opened and a silver coin is found. What is the probability that the other drawer in that cabinet contains a silver coin? The back of the book says "2/3" I cannot reason through to that answer! Ideas?

OpenStudy (anonymous):

only the drawer B and C has silver coins . so the opened drawer maybe any one of B or C. Now total no. of unopened drawer is 3, and remained silver coin is 2 as there were 4 drawers and 3 silver coins and u opened one. so the probability of getting silver coin is 2/3

OpenStudy (anonymous):

@ john farmer got it??

OpenStudy (anonymous):

if u get it then let me know.

OpenStudy (anonymous):

Isn't 2/3 the probability of getting silver if i open a random drawer from B or C? i think that the probability of getting silver by opening the other drawer in the same cabinst is .5!

OpenStudy (anonymous):

probability 5! ???

OpenStudy (anonymous):

point 5

OpenStudy (anonymous):

see the problem says u have found a silver coin, that means u have opened either B or C . do u agree??

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

i think u might wanna use bayes thm

OpenStudy (anonymous):

ok then in B and C there are total 4 drawers and in these Drawers there are total 3 silver coins. do u agree??

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

and u have opened one of them so remaining unopened drawer is 3 and the no. of remaining silver coin is 2 . do u agree??

OpenStudy (anonymous):

it asks for the prob that the OTHER drawer in the same cabinet has silver

OpenStudy (anonymous):

But the questions is not the prob of getting a silver coin by opening a random drawer You have to open the same cabinet

OpenStudy (anonymous):

yes so what??

OpenStudy (anonymous):

i think uve got to use bayes thm to show how much prob of the silver coin coming from B

OpenStudy (anonymous):

yes use it i think it should give 2/3

OpenStudy (anonymous):

@ him use bayes theorem...

OpenStudy (anonymous):

yes i got it

OpenStudy (anonymous):

[0.33 x 1]/[0.33 x 1 + 0.33x0.5]

OpenStudy (anonymous):

its 2/3 man

OpenStudy (anonymous):

this comes out to 2/3

OpenStudy (anonymous):

use fractions and ull see

OpenStudy (anonymous):

yes i know it...

OpenStudy (anonymous):

cancel 0.33 and u get 1/1.5 = 2/3

OpenStudy (anonymous):

Why are you multiplying by 1?

OpenStudy (anonymous):

bcoz in B the prob of finding a silver coin is 1

OpenStudy (anonymous):

\[P(A|B) = \frac{P(B|A) P(A)}{P(B|A) P(A) + P(B|A^{^c}) P(A^{^c})} \]he used this

OpenStudy (anonymous):

wah yaar kya mehnat ki hai

OpenStudy (anonymous):

did u get it john?

OpenStudy (anonymous):

Still puzzled.

OpenStudy (anonymous):

do u know bayes theorem?

OpenStudy (anonymous):

Yes, though I can't do it without making a tree diagram

OpenStudy (anonymous):

make a tree diagram.///thats the best way to do it

OpenStudy (anonymous):

Can you tell me where each value came from in your eq? I am not seeing the tree!

OpenStudy (anonymous):

lets take event A is for choosing box B and event B is finding 1st silver coin. so P(A|B) u want to find. got it??

OpenStudy (anonymous):

respond me...otherwise i could not understand that whether u got it or not.

OpenStudy (anonymous):

Sorry - had 2 go. Restating the question " What is prob of cabinet B give silver?" totally helped. Thanks. That gives: (Ways to get silver from B)/(Total ways to get silver) \[((1/3)(1/2) +(1/3)(1/2))/((1/3)(1/2)+(1/3)(1/2) +(1/3)(1/2))=2/3\] I still think that the way the question is phrased does not match the question " What is prob of Cabinet B given silver?" even though the author intended it to be solved using that concept. I think it is saying"What is the prob that <your next move> which has to be opening the other drawer in the same cabinet will get you silver? That probability is 50-50. Thanks again!